# Do machines fail like humans? A human-centred out-of-distribution spectrum for mapping error alignment

### 08 Mar 2026 v1

@@ 3.3 Trajectories of Alignment: Models versus humans
The trajectories of model accuracy and model–human alignment across distortion levels (Figure 4) reveal that different distortion types elicit markedly different responses from models, both in terms of accuracy and alignment with human behavior. These divergences in accuracy and alignment profiles indicate that models exploit different visual features from humans for these classification tasks. When a model is more sensitive than humans to a particular distortion, it suggests the model relies more heavily on features that this distortion selectively disrupts. Conversely, when a model is more resilient than humans, it may exploit features that humans either do not use or weight less heavily.
@@ 3.4 Alignment fingerprints across model families
As noted above, humans show roughly uniform EC with other humans for all near-OOD distortions, but model–human EC is substantially lower even for these mild distortions. This indicates that models are impaired by distortions in ways that humans are not. Moreover, different model families appear to be impaired differently by various distortions, suggesting that certain families rely on visual features that others do not, or weigh differently. Human–human MA is also relatively high for near-OOD distortions, and though lower, model–human MA remains moderately high. This implies that some false leads for humans are the same false leads for models too, particularly for certain distortion types such as power equalisation. In the far-OOD regime, EC and MA are reduced more sharply for some distortion types than others, and these patterns also differ across model families.
The radar plot (Figure 5) shows that the profile of model–human alignment across distortion types serves as a characteristic fingerprint of each model family. This fingerprint captures how a given architecture responds to different forms of image degradation relative to human perception. Our quantitative analysis reinforces this interpretation. Models are more alike within superfamilies than between superfamilies, and more alike within subfamilies than with other models in the same superfamily, although whether this reflects purely architectural effects or whether training also plays a role remains unclear. We interpret these fingerprints as evidence that model families differ in the visual features they exploit, and hence in their sensitivity to distortions that impair those features.
@@ 3.5 Model-human alignment rankings shift across the OOD spectrum
We find that VLMs were most consistently aligned with human behavior across the spectrum, CNNs were more aligned than ViTs for near-OOD, and ViTs were comparable to VLMs and better than CNNs for far-OOD regimes. These regime-dependent patterns reveal how architectural biases manifest differently under varying levels of degradation. CNNs achieve high near-OOD alignment despite their texture bias [13], which may reflect two possibilities. First, when texture and shape cues remain partially intact in near-OOD, CNNs align with humans while relying on fundamentally different features – texture and shape are often correlated in images, allowing different strategies to converge on similar outputs. Second, under moderate degradation, humans may draw on multiple features rather than relying on shape alone such as exploiting informative texture cues, thereby aligning more with CNNs. CNNs’ behavior in far-OOD catastrophically diverges from humans, with EC and MA approaching zero. This reveals that the representations CNNs develop do not align with human behavior when information becomes severely degraded, and that high alignment in normal to mildly degraded conditions does not guarantee alignment in general.
@@
VLMs demonstrate sustained strong alignment across both regimes. Notably, VLMs with CNN-based visual encoders show higher alignment in near-OOD than far-OOD, reflecting architectural biases inherited from the visual backbone. BLIP’s exceptional cross-regime stability may reflect its training approach: bootstrapping with synthetic captions may result in more robust semantic representations that generalise across degradation levels. VLMs language-based semantic knowledge may guide its behavior toward human-like decisions, especially in near-OOD. However, for far-OOD conditions, visual information may have degraded to the extent that semantic information is no longer available, leading to similar human alignment levels to the image only version of their model backbone. Language-grounded representations may guide VLMs toward human-like confusion patterns. This parallels how human semantic knowledge constrains perceptual interpretation under uncertainty [28, 42], suggesting that multimodal training provides semantic scaffolding that remains accessible even when low-level visual features become unreliable.
@@
3.6 Future Work
Future research could conduct larger-scale human studies with denser parameter sampling across distortion severity levels would strengthen conclusions about regime boundaries, better characterise the far-OOD transitional regime, and provide more reliable estimates of human–human baseline alignment. Second, our spectrum is constructed from synthetic corruptions; extending this framework to natural distributional shifts would test generalisability. Third, we focus on object recognition with static images; extending to domains such as scene understanding or video would test whether the architectural biases we observe are domain-general.
@@ 4.2 Tested Models
We evaluated 31 models spanning three architectural families: CNNs, ViTs, and VLMs. CNNs included VGG variants (VGG-11, -13, -16, -19) [35], ResNet variants (ResNet-18, -50, -101, -152) [15], DenseNet variants (DenseNet-121, -161, -169, -201) [20], Inception-v3 [37], and EfficientNet [38]. ViTs included four vanilla ViT variants [8], three Swin Transformers [27], and MaxViT-L [40]. All CNNs and most ViTs were pretrained on ImageNet-1K; EfficientNet was pretrained on a larger dataset. For VLMs, we tested BLIP [25], CLIP-ViT variants trained on DFN-2B, CLIP-ViT variants trained by OpenAI, and CLIP-ResNet variants trained by OpenAI [33]. Model details are presented in the supplementary materials.***https://docs.google.com/spreadsheets/d/1ldqG8LlQd_tDh9f3xdZQRA3v2UffUIhaLds2wTrhoLU/edit?usp=sharing

### 11 Apr 2026 v2

@@ 3.3 Trajectories of Alignment: Models versus humans
The trajectories of model accuracy and model–human alignment across distortion levels (Figure 3) reveal that different distortion types elicit markedly different responses from models, both in terms of accuracy and alignment with human behavior. These divergences indicate that models exploit different visual features to humans. When a model is more sensitive than humans to a particular distortion, it suggests that the model relies more heavily on features that this distortion disrupts. Conversely, when a model is more resilient than humans, it exploits features that humans do not use or weigh less heavily.
@@ 3.4 Alignment fingerprints across model families
As noted above, humans show roughly uniform EC with other humans for all near-OOD distortions, but model–human EC is substantially lower even for these mild distortions. This indicates that models are impaired by distortions in ways that humans are not. Moreover, different model families appear to be impaired differently by various distortions, suggesting that certain families rely on visual features that others do not, or exploit features differently. Human–human MA is also relatively high for near-OOD distortions, and though lower, model–human MA remains moderately high. This implies that some false leads for humans are the same false leads for models too – there is some overlap between the visual features exploited by machines and humans, particularly for certain distortion types such as power equalisation. In the far-OOD regime, EC and MA are reduced more sharply for some distortion types than others, and these patterns also differ across model families.
The radar plot (Figure 4) shows that the profile of model–human alignment across distortion types serves as a characteristic fingerprint of each model family. This fingerprint captures how a given architecture responds to different forms of degradation relative to human perception. Our quantitative analysis reinforces this interpretation. Models are more alike within superfamilies than between superfamilies, and more alike within subfamilies than with other models in the same superfamily, although whether this reflects architectural effects or if training also plays a role remains unclear. We interpret these fingerprints as evidence that model families differ in the visual features they exploit, and hence in their sensitivity to distortions that impair those features.
@@ 3.5 Model-human alignment rankings shift across the OOD spectrum
We find that VLMs were most consistently aligned with human behavior across the spectrum, CNNs were more aligned than ViTs for near-OOD, and ViTs were comparable to VLMs and better than CNNs for far-OOD regimes. These regime-dependent patterns reveal how architectural biases manifest differently under varying levels of degradation. CNNs achieve high near-OOD alignment despite their texture bias [13], which may reflect two possibilities. First, when texture and shape cues remain partially intact in near-OOD, CNNs align with humans while relying on fundamentally different features – one possibility is that texture and shape are often correlated in images, allowing different strategies to converge on similar outputs. Second, under moderate degradation, humans may draw on multiple features such as informative texture cues rather than relying on shape alone, thereby aligning more with CNNs. CNNs’ behavior in far-OOD catastrophically diverges from humans, with EC and MA approaching zero. This reveals that the representations CNNs develop do not align with human behavior when information is severely degraded, and that high alignment in normal to mildly degraded conditions does not guarantee alignment in general.
@@
VLMs demonstrate strong alignment across both regimes. Notably, VLMs with CNN-based visual encoders show higher alignment in near-OOD than far-OOD, reflecting architectural biases inherited from the visual backbone. BLIP’s exceptional cross-regime stability may reflect its training approach: bootstrapping with synthetic captions may result in more robust semantic representations that generalize across degradation levels. VLMs’ language-based semantic knowledge may guide its behavior toward human-like decisions, especially in near-OOD. However, for far-OOD conditions, visual information may have degraded to the extent that semantic information is no longer available, leading to similar human alignment levels to the image-only version of their model. Language-grounded representations may guide VLMs toward human-like confusion patterns. This would parallel how human semantic knowledge constrains perceptual interpretation under uncertainty [28, 42], suggesting that multimodal training provides semantic scaffolding that remains accessible even when low-level visual features become unreliable.
@@ 4.2 Tested Models
We evaluated 31 models spanning three architectural families: CNNs, ViTs, and VLMs. CNNs included VGG variants (VGG-11, -13, -16, -19) [35], ResNet variants (ResNet-18, -50, -101, -152) [15], DenseNet variants (DenseNet-121, -161, -169, -201) [20], Inception-v3 [37]. ViTs included four vanilla ViT variants [8], three Swin Transformers [27], MaxViT-L [40] and EfficientNet [38]. All CNNs and most ViTs were pretrained on ImageNet-1K; EfficientNet was pretrained on a larger dataset. For VLMs, we tested BLIP [25], CLIP-ViT variants trained on DFN-2B, CLIP-ViT variants trained by OpenAI, and CLIP-ResNet variants trained by OpenAI [33]. Model details are presented in the supplementary materials.***https://docs.google.com/spreadsheets/d/1ldqG8LlQd_tDh9f3xdZQRA3v2UffUIhaLds2wTrhoLU/edit?usp=sharing